OpenStudy (anonymous):

The area of a kite is 116.25 ft2, one diagonal measures 18.6 ft, find the measure of the other diagonal. A. 6.25 ft B. 12.5 ft C. 3.125 ft D. 25 ft

OpenStudy (campbell_st):

the area of a kite is A = 1/2 times the product of the diagonals so you know \[116.25 = \frac{1}{2} \times 18.6 \times d_{2}\] solve for the 2nd diagonal.

OpenStudy (campbell_st):

thats correct.

OpenStudy (anonymous):

B is correct ?

OpenStudy (anonymous):

hah i thought i got it wrong